OpenStudy (anonymous):

A bag of 29 tulip bulbs contains 10 red tulip bulbs, 11 yellow tulip bulbs, and 8 purple tulip bulbs. What is the probability that two randomly selected tulip bulbs are both red? What is the probability that the first bulb selected is red and the second is yellow? What is the probability that the first bulb selected is yellow and the second is red?

OpenStudy (anonymous):

@nechirwan

OpenStudy (anonymous):

@sasogeek

OpenStudy (kendricklamar2014):

Probability first bulb is red = 14/29 There are not 13 red bulbs left and 28 total bulbs left Probability second bulb is red = 13/28 p = 14/29 * 13/28 = 13/58 = 0.224 2) Probability first bulb is red = 14/29 Probability second bulb is yellow = 8/28 p = 14/29 * 8/28 = 4/29 = 0.138 3) Probability first bulb is yellow = 8/29 Probability second bulb is red = 14/28 p = 8/29 * 14/28 = 4/29 = 0.138

OpenStudy (sasogeek):

@KendrickLamar2014 where's the 14/29 from?

OpenStudy (kendricklamar2014):

14+8+7 = 29

OpenStudy (sasogeek):

it's 10 red, 11 yellow and 8 purple... did you read the question?

OpenStudy (anonymous):

I need the probability for both bulbs being red :/

OpenStudy (kendricklamar2014):

Whoops, OS was lagging

OpenStudy (sasogeek):

probability of first bulb being red is 10/29... if that one is taken out, you have 28 bulbs left total and 9 red bulbs left, probability of second ball being red is 9/28. since it's first AND second bulb being red, you have 10/29 * 9/28 = probability of both being red

OpenStudy (sasogeek):

bulb*

OpenStudy (sasogeek):

do you understand?